# Measuring multipartite quantum correlations by thermodynamic work extraction

### v1

@@ Appendix A Relationship to other measures of multipartite quantum correlations
We discuss the relationship between the multipartite work deficit introduced in this work and other existing measures of multipartite quantum correlations, especially in terms of the numerical cost. Many existing measures, such as the relative entropy of entanglement Plenio and Vedral (2001); Wei et al. (2004), geometric entanglement Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995), and global quantum discord Rulli and Sarandy (2011), involve optimization parameters proportional to the system size, similarly to the multipartite work deficit. This increase of the optimization parameters makes it difficult to calculate these measures in many-body systems  [^3] We note that in some previous works Orús (2008); Shi et al. (2010); Orús and Wei (2010); Mousolou et al. (2013); Huang and Lin (2010); Orus and Wei (2010); Sun et al. (2015); Huang et al. (2018), they avoid this cost increase by posing some assumptions on the optimal parameters.. While our efficient calculation method can reduce the numerical cost for the multipartite work deficit and global quantum discord  [^4] See Supplemental Material for the details., extending this method to other measures remains as a future issue.
@@ B.1 Expression of $W_{l}$
We here derive the expression for $W_{l}$ in Eq. (4). For this purpose, we show that for any measurement protocol $\Lambda$, which includes higher-rank projectors, there exists a better protocol $\Lambda^{\prime}$ consisting solely of rank-one projective measurements, where $W_{\Lambda}\leq W_{\Lambda^{\prime}}$ holds.
By denoting the local projectors in protocol $\Lambda$ as $\{\Pi_{y_{1}}\},\{\Pi_{y_{2}}\},\dots,\{\Pi_{y_{N}}\}$, the conditional density operator is defined as
@@
where $p_{\Lambda}(y_{n})$ represents the probability of the measurement outcome $y_{n}$. The measurement protocol $\Lambda^{\prime}$, which satisfies $W_{\Lambda}\leq W_{\Lambda^{\prime}}$, can be constructed by replacing each higher-rank projector $\Pi_{y_{n}}$ in $\Lambda$ with its fine-grained rank-one projectors $\{\Pi_{x_{n}}^{y_{n}}\Pi_{y_{n}}\}_{x_{n}}$, where $\Pi_{x_{n}}^{y_{n}}$ denotes the diagonal bases of the conditional density operator $\widetilde{\rho}_{n}^{\Lambda,y_{n}}=\sum_{x_{n}}q^{y_{n}}(x_{n})\Pi_{x_{n}}^{y_{n}}$. By denoting the measurement outcomes for the new protocol $\Lambda^{\prime}$ as $(Y_{n},X_{n})\equiv(y_{1},x_{1},y_{2},x_{2},\dots,y_{n},x_{n})$, we can show the desired inequality as follows:
@@ B.2 Expression with quantum relative entropy
We here show the equivalence of the two expressions of multipartite work deficit, namely Eq. (5) and Eq. (6). Equation (5) can be transformed into the representation with quantum relative entropy as
$\begin{split}\min_{\Lambda\in\mathcal{M}_{N}}H_{\Lambda}(Y_{N})-S(\rho)&=\min_{\Lambda\in\mathcal{M}_{N}}S\left(\pi_{\Lambda}(\rho)\right)-S(\rho)\\ &=\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\pi_{\Lambda}(\rho)\right),\end{split}$ (49)
where $\pi_{\Lambda}$ denotes the dephasing map with respect to the measurement basis of $\Lambda$. Since $\pi_{\Lambda}(\rho)\in\mathcal{C}_{N}$, the following inequality holds:
$\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\pi_{\Lambda}(\rho)\right)\geq\min_{\sigma\in\mathcal{C}_{N}}S(\rho\|\sigma).$ (50)
@@
$S(\rho\|\sigma)=S\left(\rho\|\pi_{\sigma}(\rho)\right)+S\left(\pi_{\sigma}(\rho)\|\sigma\right),$
where $\pi_{\sigma}$ denotes the dephasing map with respect to the diagonal basis of $\sigma$. Therefore, we can derive the following inequality:
$\begin{split}\min_{\sigma\in\mathcal{C}_{N}}S(\rho\|\sigma)&=\min_{\sigma\in\mathcal{C}_{N}}S\left(\rho\|\pi_{\sigma}(\rho)\right)+S\left(\pi_{\sigma}(\rho)\|\sigma\right)\\ &\geq\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\pi_{\Lambda}(\rho)\right).\end{split}$ (51)
From Eqs. (49), (50) and (51) we can derive the equivalence of Eq. (5) and Eq. (6) as follows:
@@ Appendix C Relationship between multipartite work deficit and bipartite quantum correlation
$\Delta\geq\sum_{n=1}^{N-1}D_{(1,\dots,n);n+1}.$ (11)
which shows the relationship between multipartite work deficit and the well-known measure of bipartite quantum correlation $D_{(1,\dots,n);n+1}$ called the quantum discord. The quantum discord for the bipartite state $\rho_{1,2}$ is defined as
@@
where $\mathcal{M}_{1}$ is the set of projective measurement on party $a_{1}$, $p_{\Lambda}(y_{1})$ is the probability of obtaining measurement outcome $y_{1}$, $\rho_{2}^{\Lambda,y_{1}}$ is the conditional density operator on party $a_{2}$ given the outcome $y_{1}$, and $S_{2|1}(\rho_{1,2})\equiv S(\rho_{1,2})-S(\rho_{1})$ denotes the conditional von Neumann entropy. The inequality (11) can be derived as follows:
@@ Appendix D Efficient calculation method for work deficit density $\delta$
D.1 Lower bounds
We provide detailed explanation on the lower bound $L_{l}$. The lower bound is defined as
where $\mathcal{N}_{n:l}$ is the set of measurement protocol whose POVM operators are represented as
$\{E_{Z_{n}}\otimes\Pi_{y_{n+1}}\otimes\Pi_{y_{n+2}}\dots\otimes\Pi_{y_{n+l}}\}_{Z_{n},Y_{n:l}}.$
Here, $\{E_{Z_{n}}\}$ are the operators for the arbitrary POVM on parties $\{a_{i}\}_{i=1}^{n}$, whose Hilbert space dimension become $d^{n}$. Due to this exponential growth of the system dimension, it seems very difficult to perform optimization over the arbitrary POVM at first glance. However, as mentioned in the main text, the optimization region can be greatly reduced when the state is described with MPS with fixed bond dimension $D_{B}$. To be specific, we prove that optimal protocol can be found in the set of rank-one extremal POVM on the support of $\rho_{1,\dots,n}$ denoted as $\mathrm{supp}[\rho_{1,\dots,n}]$, as shown in Fig. 12. For an MPS with a bond dimension $D_{B}$, the dimension of $\mathrm{supp}[\rho_{1,\dots,n}]$ is also $D_{B}$. Therefore, it is only necessary to consider rank-one extremal POVMs on a $D_{B}$-dimensional system. Since such POVMs can be represented with fewer than $D_{B}^{2}$ operators D’Ariano et al. (2005), this optimization can be performed with a constant numerical cost that is independent of the system size.
Figure 12: Optimization region for the POVM on $\{a_{i}\}_{i=1}^{n}$ for the lower bounds $L_{l}$ and $L_{l}^{\to}$. Although the lower bounds $L_{l}$ and $L_{l}^{\to}$ require optimization over arbitrary POVMs on the parties $\{a_{i}\}_{i=1}^{n}$, this optimization region can be reduced to extremal rank-one POVMs on the support of $\rho_{1,\dots,n}$. This reduction is particularly effective for states well-approximated by MPS.
The reduction of the optimization region illustrated in Fig. 12 is valid only for states $\rho$ with a fixed bond dimension independent of $n$. However, we can show that this cost reduction is also useful for the state $\rho$ well-approximated by a state $\rho^{\prime}$ with the fixed bond dimension $D_{B}$. Specifically, we derive that
$|L_{l}(\rho)-L_{l}(\rho^{\prime})|\leq\tilde{\mathcal{O}}(\mu)$ (55)
holds for states $\rho$ and $\rho^{\prime}$ satisfying $\|\rho-\rho^{\prime}\|_{\rm tr}=\mu$, where $\|X\|_{\rm tr}\equiv\mathrm{Tr}[|X|]$ denotes the trace norm. Here, $L_{l}(\rho)$ represents the lower bound for the state $\rho$. Consequently, if the state $\rho$ is well-approximated by the state $\rho^{\prime}$ with a fixed bond dimension, the work deficit density is lower-bounded as
$\delta(\rho)\geq L_{l}(\rho^{\prime})-\tilde{\mathcal{O}}(\mu),$
where $L_{l}(\rho^{\prime})$ is efficiently calculable. The detailed proof of the reduction of the optimization region shown in Fig. 12 and the continuity of the lower bound $L_{l}$ in Eq. (55) is provided in Supplemental Material.
D.2 Upper bounds
We provide detailed explanation on the upper bounds $U_{l}$ and $U_{l,k}$. The upper bounds $U_{l}$ is defined as
@@ D.2 Upper bounds
$\{\Pi_{y_{n+1}}\otimes\Pi_{y_{n+2}}\dots\otimes\Pi_{y_{n+l}}\}_{Y_{n:l}},$
where $\{\Pi_{y_{n+i}}\}$ represents the local projective measurement on party $a_{n+i}$. Especially, for the translationally invariant system under periodic boundary condition, the upper bound $U_{l}$ is reduced to
@@
since all the terms $\{H_{\Lambda}(Y_{n:l})\}_{n\in I_{l}}$ are identical due to the translational invariance.
For such states, we can further introduce another type of upper bound $U_{l,k}$. It is defined as
@@
where $\mathcal{M}_{l}$ is the set of the measurement protocols on the $l$-site subsystem, and the measurement protocol $\Lambda^{\otimes(k+1)}$ is to perform the identical measurement $\Lambda$ on $(k+1)$ sequential $l$-site segments, specifically $\{a_{i}\}_{i=n-kl+1}^{n-kl+l},\dots,\{a_{i}\}_{i=n+1}^{n+l}$. We note that in the infinite systems, the first term in Eq. (57) is negligible. By comparing Eq. (56) to Eq. (57), we can see that the upper bound $U_{l,0}$ is reduced to $U_{l}$, and therefore $U_{l,k}$ is the generalization of $U_{l}$. We can derive the upper bound $U_{l,k}$ as follows:
$\delta$ $=\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{N}}\left\{\sum_{n\in I_{l}}H_{\Lambda}(Y_{n:l}|Y_{n})\right\}$
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{N}}\left\{\sum_{n\in I_{l}}H_{\Lambda}(Y_{n:l}|Y_{n-kl:kl})\right\}$
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{l}}\left\{\sum_{n\in I_{l}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\right\}$
@@
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{l}}\left\{kl\ln d+\sum_{\begin{subarray}{c}n\in I_{l}\\ n\geq kl\end{subarray}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\right\}$
$=\frac{kl}{N}\ln d+\frac{1}{N}\left(\frac{N}{l}-k\right)\min_{\Lambda\in\mathcal{M}_{l}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})$
$\leq U_{l,k},$
@@
where we here assume $S(\rho)=0$ while the derivation for the case of $S(\rho)\neq 0$ follows in the same way. The equality in the fourth line follows from the fact that the terms $\{H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\}_{n\in I_{l},n\geq kl}$ are identical due to the translational invariance.
D.3 Tightness of the lower and upper bounds

### v2

@@ Appendix A Relationship to other measures of multipartite quantum correlations
We discuss the relationship between the multipartite work deficit introduced in this work and other existing measures of multipartite quantum correlations, especially in terms of the numerical cost. Many existing measures, such as the relative entropy of entanglement Plenio and Vedral (2001); Wei et al. (2004), geometric entanglement Barnum and Linden (2001); Wei and Goldbart (2003); Shimony (1995), and global quantum discord Rulli and Sarandy (2011), involve optimization parameters proportional to the system size, similarly to the multipartite work deficit. This increase of the optimization parameters makes it difficult to calculate these measures in many-body systems. While our efficient calculation method can reduce the numerical cost for the multipartite work deficit and global quantum discord, extending this method to other measures remains as a future issue.
@@ B.1 Expression of $W_{l}$
We here show that extractable work through zero-way closed LOCC is expressed as
$W_{l}=N\ln d-\min_{\Lambda\in\mathcal{M}_{N}}H_{\Lambda}(Y_{N}).$ (4)
To derive this expression from Eq. (3) in the main text, we need to show that optimal measurement protocol is always rank-one. In the following, we prove that for any measurement protocol $\Lambda$ which includes higher-rank projectors, there exists a certain protocol $\Lambda^{\prime}$ consisting solely of rank-one projectors, such that $W_{\Lambda^{\prime}}\geq W_{\Lambda}$ is satisfied.
In the zero-way closed LOCC, the allowed local measurement on each party is not arbitrary POVM but projective measurement. Therefore, the measurement protocol $\Lambda$ is described with the sets of orthogonal projectors, represented as $\{\Pi_{y_{1}}\},\{\Pi_{y_{2}}\},\dots,\{\Pi_{y_{N}}\}$. In this case, the conditional density operator on the $n$-th party is defined as
@@
where $p_{\Lambda}(y_{n})$ represents the probability of the measurement outcome $y_{n}$. The rank-one measurement protocol $\Lambda^{\prime}$, which satisfies $W_{\Lambda^{\prime}}\geq W_{\Lambda}$, can be constructed by replacing each higher-rank projector $\Pi_{y_{n}}$ in $\Lambda$ with its fine-grained rank-one projectors $\{\Pi_{x_{n}}^{y_{n}}\Pi_{y_{n}}\}_{x_{n}}$, where $\{\Pi_{x_{n}}^{y_{n}}\}_{x_{n}}$ is the projectors on the diagonal basis of the conditional density operator defined as $\widetilde{\rho}_{n}^{\Lambda,y_{n}}=\sum_{x_{n}}q^{y_{n}}(x_{n})\Pi_{x_{n}}^{y_{n}}$. By representing the measurement outcomes for the new protocol $\Lambda^{\prime}$ as $(Y_{n},X_{n})\equiv(y_{1},x_{1},y_{2},x_{2},\dots,y_{n},x_{n})$, we can derive the inequality $W_{\Lambda^{\prime}}\geq W_{\Lambda}$ as follows:
@@ B.2 Expression with quantum relative entropy
We here show the equivalence of the two expressions of multipartite work deficit, namely Eqs. (5) and (6). Equation (5) can be transformed into the representation with quantum relative entropy as
$\begin{split}\min_{\Lambda\in\mathcal{M}_{N}}H_{\Lambda}(Y_{N})-S(\rho)&=\min_{\Lambda\in\mathcal{M}_{N}}S\left(\mathcal{D}_{\Lambda}(\rho)\right)-S(\rho)\\ &=\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\mathcal{D}_{\Lambda}(\rho)\right),\end{split}$ (49)
where $\mathcal{D}_{\Lambda}$ denotes the dephasing map with respect to the measurement basis of $\Lambda$. Since $\mathcal{D}_{\Lambda}(\rho)\in\mathcal{C}_{N}$, the following inequality holds:
$\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\mathcal{D}_{\Lambda}(\rho)\right)\geq\min_{\sigma\in\mathcal{C}_{N}}S(\rho\|\sigma).$ (50)
@@
$S(\rho\|\sigma)=S\left(\rho\|\mathcal{D}_{\sigma}(\rho)\right)+S\left(\mathcal{D}_{\sigma}(\rho)\|\sigma\right),$
where $\mathcal{D}_{\sigma}$ denotes the dephasing map with respect to the diagonal basis of $\sigma$. Therefore, we can derive the following inequality:
$\begin{split}\min_{\sigma\in\mathcal{C}_{N}}S(\rho\|\sigma)&=\min_{\sigma\in\mathcal{C}_{N}}S\left(\rho\|\mathcal{D}_{\sigma}(\rho)\right)+S\left(\mathcal{D}_{\sigma}(\rho)\|\sigma\right)\\ &\geq\min_{\Lambda\in\mathcal{M}_{N}}S\left(\rho\|\mathcal{D}_{\Lambda}(\rho)\right).\end{split}$ (51)
From Eqs. (49), (50), and (51), we can derive the equivalence of Eqs. (5) and (6) as follows:
@@ Appendix C Relationship between multipartite work deficit and bipartite quantum correlation
$\Delta\geq\sum_{n=1}^{N-1}D_{(1,\dots,n);n+1},$ (11)
which indicates the relationship between multipartite work deficit and quantum discord $D_{(1,\dots,n);n+1}$, a well-known measure of bipartite quantum correlation Ollivier and Zurek (2001); Henderson and Vedral (2001). The quantum discord for a bipartite state $\rho_{1,2}$ is defined as
@@
where $\mathcal{M}_{1}$ is the set of projective measurements on party $a_{1}$, $p_{\Lambda}(y_{1})$ is the probability of obtaining measurement outcome $y_{1}$, $\rho_{2}^{\Lambda,y_{1}}$ is the conditional density operator on party $a_{2}$ given the outcome $y_{1}$, and $S_{2|1}(\rho_{1,2})\equiv S(\rho_{1,2})-S(\rho_{1})$ denotes the conditional von Neumann entropy. The inequality (11) can be derived as follows:
@@ Appendix D Efficient calculation method for work deficit density $\delta$
D.1 Upper bound $U_{l,k}$
We here introduce the definition of the upper bound $U_{l,k}$ of work deficit density $\delta$, which is defined for the translationally invariant state with periodic boundary condition. This is regarded as a generalization of the upper bound $U_{l}$, which is defined as follows:
@@ D.1 Upper bound $U_{l,k}$
$\{\Pi_{y_{n+1}}\otimes\Pi_{y_{n+2}}\dots\otimes\Pi_{y_{n+l}}\}_{Y_{n:l}}.$
Here, $\{\Pi_{y_{n+i}}\}$ represents the set of orthogonal rank-one projectors on party $a_{n+i}$. For the translationally invariant system under periodic boundary condition, $U_{l}$ is represented as
@@
since all the terms in the summation $\left\{\min_{\Lambda\in\mathcal{M}_{n:l}}H_{\Lambda}(Y_{n:l})\right\}_{n\in I_{l}}$ are identical.
For such states, the upper bound $U_{l,k}$ is defined as follows, which is reduced to Eq. (55) when $k=0$:
@@
where $\mathcal{M}_{l}$ is the set of local projective measurements on the $l$-site segment, and the measurement protocol $\Lambda^{\otimes(k+1)}$ is to perform the identical measurement $\Lambda$ on $(k+1)$ sequential $l$-site segments, specifically $\{a_{i}\}_{i=n-kl+1}^{n-kl+l},\dots,\{a_{i}\}_{i=n+1}^{n+l}$. We note that in the thermodynamic limit $N\to\infty$, the first term in Eq. (56) is negligible. We can derive the inequality $\delta\leq U_{l,k}$ as follows:
$\delta$ $=\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{N}}\left\{\sum_{n\in I_{l}}H_{\Lambda}(Y_{n:l}|Y_{n})\right\}-\frac{S(\rho)}{N}$
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{N}}\left\{\sum_{n\in I_{l}}H_{\Lambda}(Y_{n:l}|Y_{n-kl:kl})\right\}-\frac{S(\rho)}{N}$
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{l}}\left\{\sum_{n\in I_{l}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\right\}-\frac{S(\rho)}{N}$
@@
$\leq\frac{1}{N}\min_{\Lambda\in\mathcal{M}_{l}}\left\{\sum_{\begin{subarray}{c}n\in I_{l}\\ n\geq kl\end{subarray}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\right\}$
$\quad\quad\quad\quad\quad\quad+\frac{kl\ln d}{N}-\frac{S(\rho)}{N}$
$=\left(\frac{1}{l}-\frac{k}{N}\right)\min_{\Lambda\in\mathcal{M}_{l}}H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})$
$\quad\quad\quad\quad\quad\quad+\frac{kl\ln d}{N}-\frac{S(\rho)}{N}$
$\leq U_{l,k}.$
@@
The equality in the fifth line follows from the fact that all the terms in the summation $\{H_{\Lambda^{\otimes(k+1)}}(Y_{n:l}|Y_{n-kl:kl})\}_{n\in I_{l},n\geq kl}$ are identical due to the translation invariance of the system and measurement protocol.
D.2 Reduction of the optimization region for the lower bounds
Figure 12: Reduction of the optimization region in the calculation of lower bound $L_{l}$. Although the lower bound $L_{l}$ includes the optimization over the set of arbitrary POVMs on the parties $\{a_{i}\}_{i=1}^{n}$, denoted as $\mathcal{E}_{n}$, this optimization region can be effectively reduced to the set of extremal rank-one POVMs on $\mathrm{Row}[\rho_{1,\dots,n}]$, denoted as $\mathcal{E}^{\prime}_{n}$ (Theorem 1). Especially for the MPS with fixed bond dimension $D_{B}$, optimization over the reduced region $\mathcal{E}^{\prime}_{n}$ can be performed at a constant cost independent of $n$.
We provide a detailed discussion on the optimization involved in the definition of the lower bound $L_{l}$. The lower bound is defined as
$L_{l}\equiv\frac{1}{N}\left\{\sum_{n\in I_{l}}\min_{\mu\in\mathcal{E}_{n}}\min_{\pi\in\mathcal{M}_{n:l}}H_{(\mu,\pi)}(Y_{n:l}|Z_{n})-S(\rho)\right\},$ (57)
where $\mathcal{E}_{n}$ denotes the set of arbitrary POVMs on the parties $\{a_{i}\}_{i=1}^{n}$, and $\mathcal{M}_{n:l}$ represents the set of local projective measurements on the neighbouring $l$ parties $\{a_{i}\}_{i=n+1}^{n+l}$. Equation (57) is equivalent to the representation in Eq. (19), since the measurement protocol $\Lambda\in\mathcal{N}_{n:l}$ is divided into $\mu\in\mathcal{E}_{n}$ and $\pi\in\mathcal{M}_{n:l}$ as $\Lambda=(\mu,\pi)$. The measurement operators for $(\mu,\pi)$ are given by
$\{E_{Z_{n}}\otimes\Pi_{y_{n+1}}\otimes\Pi_{y_{n+2}}\dots\otimes\Pi_{y_{n+l}}\}_{Z_{n},Y_{n:l}},$
where $\{E_{Z_{n}}\}$ is the set of POVM operators for $\mu$, and $\{\Pi_{y_{n+i}}\}$ is the set of orthogonal projectors on party $a_{n+i}$. Since the Hilbert space dimension of the subsystem $\{a_{i}\}_{i=1}^{n}$ is $d^{n}$, the optimization cost for $\mu\in\mathcal{E}_{n}$ appears to grow exponentially with respect to $n$, at first glance.
However, we here show that this cost can be reduced to constant independent of $n$, for MPS with fixed bond dimension $D_{B}$. While we have only provided an intuitive explanation for this cost reduction in the main text, suggesting that the subsystem $\{a_{i}\}_{i=1}^{n}$ is regarded as a fixed-dimensional system, we present a more detailed discussion in this subsection. Specifically, we show that the optimization region $\mathcal{E}_{n}$ for the POVM $\mu$ can be reduced to the set of rank-one extremal POVMs on $\mathrm{Row}[\rho_{1,\dots,n}]$ denoted as $\mathcal{E}_{n}^{\prime}$, as illustrated in Fig. 12. Here, $\mathrm{Row}[\rho_{1,\dots,n}]$ is the row space of the reduced density operator $\rho_{1,\dots,n}$, namely, the subspace spanned by the row vectors of $\rho_{1,\dots,n}$. For an MPS with bond dimension $D_{B}$, the dimension for $\mathrm{Row}[\rho_{1,\dots,n}]$ is fixed ($D_{B}$ for open boundary condition, $D_{B}^{2}$ for periodic boundary condition). A rank-one POVM is a class of POVM defined by rank-one operators $\{E_{Z_{n}}\}$, and an extremal POVM is one that cannot be described as a convex combination of other POVMs D’Ariano et al. (2005); Haapasalo et al. (2012). This reduction of optimization region is stated as the following theorem:
Theorem 1.
Let $\mathcal{E}_{n}$ be the set of arbitrary POVMs on parties $\{a_{i}\}_{i=1}^{n}$, and $\mathcal{E}_{n}^{\prime}$ be the set of rank-one extremal POVMs on $\mathrm{Row}[\rho_{1,\dots,n}]$. Then, the following equality is satisfied:
$\begin{split}L_{l}&\equiv\frac{1}{N}\left\{\sum_{n\in I_{l}}\min_{\mu\in\mathcal{E}_{n}}\min_{\pi\in\mathcal{M}_{n:l}}H_{(\mu,\pi)}(Y_{n:l}|Z_{n})-S(\rho)\right\}\\ &=\frac{1}{N}\left\{\sum_{n\in I_{l}}\min_{\mu\in\mathcal{E}_{n}^{\prime}}\min_{\pi\in\mathcal{M}_{n:l}}H_{(\mu,\pi)}(Y_{n:l}|Z_{n})-S(\rho)\right\}.\end{split}$ (58)
The proof of this theorem is provided in the Supplemental Material SM . Since the rank-one extremal POVM on $x$-dimensional system is described with fewer than $x^{2}$ rank-one measurement operators $\{E_{Z_{n}}\}$ D’Ariano et al. (2005); Haapasalo et al. (2012), the optimization of $\mu$ within $\mathcal{E}_{n}^{\prime}$ can be performed at a constant cost independent of $n$. This is crucial for the lower bound $L_{l}$ to be calculated with linear cost with respect to the system size $N$ (Table 1).
D.3 Lower and upper bounds for states approximated by MPS
While we have primarily focused on systems exactly represented by MPS, we here explain that our efficient calculation method is applicable also for states approximated by MPS. Specifically, we consider the state $\rho$ which is approximated with an MPS $\rho^{\prime}$ with fixed bond dimension. To evaluate the lower bound $L_{l}(\rho)$, we utilize the fact that it is well-approximated with $L_{l}(\rho^{\prime})$, which is efficiently computable as the consequence of Theorem 1. To be more precise, we derive the following theorem:
Theorem 2.
Let $\rho$ and $\rho^{\prime}$ be the states satisfying $\|\rho-\rho^{\prime}\|_{\rm tr}\leq\nu$, where $\|X\|_{\rm tr}\equiv\mathrm{Tr}[|X|]$ denotes the trace norm. Then, the following inequalities are satisfied:
$|U_{l}(\rho)-U_{l}(\rho^{\prime})|$ $\leq\tilde{\mathcal{O}}(\nu),$ (59)
$|L_{l}(\rho)-L_{l}(\rho^{\prime})|$ $\leq\tilde{\mathcal{O}}(\nu).$ (60)
The proof of this theorem is provided in the Supplemental Material SM . Since Eq. (60) implies the inequality
$\delta(\rho)\geq L_{l}(\rho)\geq L_{l}(\rho^{\prime})-\tilde{\mathcal{O}}(\nu),$
we can obtain the efficiently computable lower bound of $\delta(\rho)$ through the calculation in the MPS $\rho^{\prime}$.
D.4 Tightness of the lower and upper bounds